OpenStudy (anonymous):

Remainder when 23^101 is divided by 9 is?

OpenStudy (anonymous):

@experimentX @Arnab09 @yash2651995

OpenStudy (lgbasallote):

weird...my calculator cant solve 23^101

OpenStudy (anonymous):

Some tricks to solve it!!

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@Arnab09 u r correct can u show it

OpenStudy (anonymous):

u know the method or just testing? :)

OpenStudy (anonymous):

no i dont know!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

23=5(mod 9)

OpenStudy (anonymous):

so, 23^3=125(mod 9)=-1 (mod 9) so, 23^101=23^99* 529= (-1)^3*529( mod 9)= -529(mod 9)=2( mod 9)

OpenStudy (anonymous):

sorry, thats (-1)^33 ^^^up there

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

No @Arnab09 i did nt understand!

OpenStudy (experimentx):

find some cyclic pattern... try dividing few powers of 23 by 9 until it gets repeated

OpenStudy (anonymous):

23=5(mod 9) ?????????????

OpenStudy (anonymous):

u know the divisibility formulae? @Yahoo! ?

OpenStudy (anonymous):

No....CAn u show that

OpenStudy (anonymous):

that means 23 leaves remainder 5 when divided by 9

OpenStudy (experimentx):

Oh .. yes that's right!! it's equivalent to dividing 5^101 by 9

OpenStudy (anonymous):

@Arnab09 ok then this so, 23^3=125(mod 9)=-1 (mod 9) so, 23^101=23^99* 529= (-1)^3*529( mod 9)= -529(mod 9)=2( mod 9)????????

OpenStudy (anonymous):

thats (-1)^33, i told before

OpenStudy (anonymous):

it is bit confusing can u write it lol!!

OpenStudy (anonymous):

@experimentX wat is that pattern

OpenStudy (anonymous):

23^99=(23^3)^33=(-1)^33 (mod 9)=-1(mod 9) now?

OpenStudy (experimentx):

5^1 mod 9 = 5 5^2 mod 9 = 7 5^3 mod 9 = 8 5^4 mod 9 = 4 5^5 mod 9 = 2 5^6 mod 9 = 1 5^7 mod 9 = 5 <<<--- it will just repeat after that ... 7,8,4,2,1,5 ... so on and on .. so on and on ..

OpenStudy (anonymous):

as 23^3=(-1) (mod 9) as stated before..

OpenStudy (anonymous):

when we divide 23^3 by 9 we get remainder as 8 not 1

OpenStudy (anonymous):

-1 means 8 :)

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

difficult to explain..... number theory.. divisibility rule

OpenStudy (anonymous):

@experimentX if the pattern is 7,8,4,2,1,5 . then....

OpenStudy (anonymous):

find the 101 th term

OpenStudy (anonymous):

it's 2 i solve with calc

OpenStudy (anonymous):

how @Arnab09 it is not a AP so??

OpenStudy (anonymous):

the remainder is repeated every 6th time

OpenStudy (anonymous):

so, to get 101th term, it will complete 16 cycles+the 5th term.. that is 2

OpenStudy (anonymous):

ok got it thxx

OpenStudy (anonymous):

better if u apply divisibility rule.. get acquainted with it in number theory!!

OpenStudy (anonymous):

but can u give some basics abt that

OpenStudy (experimentx):

it repeats after 6th term, so 101 can be written as 16*6+5 so it reduces to 5^(16*6+5) mod 9 = 5^5 mod 9 http://www.wolframalpha.com/input/?i=5^%286*16%2B5%29+mod+9 http://www.wolframalpha.com/input/?i=23^101+mod+9 http://www.wolframalpha.com/input/?i=5^5+mod+9

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Divisibility

OpenStudy (anonymous):

u can get something here^^ @Yahoo!

OpenStudy (anonymous):

@Arnab09 look my new quest

OpenStudy (experimentx):

oh .. new trick on wolfram!!

OpenStudy (experimentx):

what is it?? LOL http://www.wolframalpha.com/input/?i=quotient+remainder+%285%5E101%2C+9%29